# Functional Central Limit Theorem for the simultaneous subgraph count of dynamic Erdős-Rényi random graphs

### v1

@@
Abstract.
In this paper we consider a dynamic Erdős-Rényi random graph with independent identically distributed edge processes. Our aim is to describe the joint evolution of the entries of a subgraph count vector. The main result of this paper is a functional central limit theorem: we establish, under an appropriate centering and scaling, the joint functional convergence of the vector of subgraph counts to a specific multidimensional Gaussian process. The result holds under mild assumptions on the edge processes, most notably a Lipschitz-type condition.
Keywords. Dynamic random graphs $\circ$ subgraph count $\circ$ functional central limit theorem
Affiliations. RH and MM are with the Mathematical Institute, Leiden University, P.O. Box 9512, 2300 RA Leiden, The Netherlands. MM is also affiliated with Korteweg-de Vries Institute for Mathematics, University of Amsterdam, Amsterdam, The Netherlands; Eurandom, Eindhoven University of Technology, Eindhoven, The Netherlands; Amsterdam Business School, Faculty of Economics and Business, University of Amsterdam, Amsterdam, The Netherlands.
NK is with the Korteweg-de Vries Institute for Mathematics, University of Amsterdam, Science Park 904, 1098 XH Amsterdam, The Netherlands.
Acknowledgments. This research was supported by the European Union’s Horizon 2020 research and innovation programme under the Marie Sklodowska-Curie grant agreement no. 945045, and by the NWO Gravitation project NETWORKS under grant agreement no. 024.002.003. [Uncaptioned image]
1. Introduction
Over the past decades, inspired by applications in a broad range of disciplines (such as biology, physics, economics, and computer science), various types of random graph models have been introduced and analyzed. Random graph models are specified via an underlying probabilistic mechanism. The Erdős-Rényi random graph [7, 11], in which there are $N$ vertices and in which every possible vertex pair is connected independently with a given probability $p\in(0,1)$, is arguably the most well-known.
Most literature on random graphs focuses on static models, which represent a single instance of the random graph without accounting for them changing over time. However, real-world networks typically do evolve over time, motivating the study of dynamic random graphs, particularly dynamic Erdős-Rényi random graphs (e.g., [25, 20, 18, 17, 5]). In these models, edges evolve independently over time, alternating between being present and absent. Essentially all properties that have been studied for static random graphs can be analyzed for their dynamic counterparts as well. For instance, in the context of dynamic random graphs there is now a notion of a giant component process, and a notion of a spectrum process. In the present paper we concentrate on the subgraph counts process, in that our aim is to describe the joint evolution of the entries of a subgraph count vector. We thus capture, for instance, the joint distribution of the number of edges and the number of triangles over time.
@@ 1. Introduction
The first results showing asymptotic normality for subgraph counts in a static Erdős-Rényi random graph appeared already in [6]: an appropriately centered and normalized version of the subgraph count converges, as $N\to\infty$, to a zero-mean Gaussian random variable. These results were further generalized to the regime where the edge probability $p=p_{N}$ tends to zero as the number of vertices $N$ grows large; see e.g. [23] that covers specific types of subgraphs. Finally, in [24] necessary and sufficient conditions on the decay of $p_{N}$ were established for obtaining asymptotic normality for any subgraph count.
@@
Importantly, there seem to be no asymptotic normality results covering the setting of random graphs that evolve over time, even in the simplest case in which the edge processes do not depend on the number of vertices $N$. The main objective of this paper is to settle this, by establishing a joint functional central limit theorem for dynamic Erdős-Rényi random graphs. Concretely, we identify conditions under which an appropriately centered and normalized version of a vector of subgraph counts has, at the process level, a Gaussian limit. Moreover, we explicitly characterize the limiting Gaussian process. The type of convergence considered is in distribution, in $\mathbb{D}([0,T],\mathbb{R}^{m})$, for some finite, deterministic horizon $T$.
In our setup, we do not focus on one specific type of edge dynamics, in that we aim to work with a specification of the dynamic Erdős-Rényi random graph that is as general as possible. We follow the ‘traditional’ path in proving process-level convergence: we first establish finite-dimensional convergence, followed by a tightness step that yields the desired functional convergence. In order to accommodate this tightness, we have to impose some assumptions on the edge dynamics. More concretely, in our framework the probabilities of the edge switching once or twice (in an interval of given length, that is) are assumed to obey a Lipschitz-type condition. At the methodological level, some of the steps of our analysis follow their counterparts in [13], where the principal eigenvalue process of the graph’s adjacency matrix is analyzed in the regime in which the edge processes do not depend on $N$.
The main result of this paper, Theorem 4.1, establishes that the limiting behavior (as the size of our graph grows large) of the subgraph count process is Gaussian, and strictly correlated with the subgraph count for the least common subgraph of our graph. This least common subgraph is defined as a subgraph with at least one edge and the asymptotically smallest expected number in our dynamic Erdős-Rényi random graph. For instance, if the edge process does not depend on the size of the random graph (as before denoted by $N$), the expected subgraph count asymptotically scales as $N^{v}$, where $v$ is the size of the subgraph under consideration. In this scenario, the least common subgraph for any graph with at least one edge is a single edge, and thus every subgraph count process is strictly correlated with the edge count. This implies that all possible subgraph counts are strictly correlated with one another. We provide a detailed exploration of this phenomenon in Corollary 4.3.
@@
2. Model
We start by constructing our dynamic Erdős-Rényi random graph process, in the sequel denoted by $G_{N}(t)$, with $N\in\mathbb{N}$ vertices over the time interval $t\in[0,T]$ for some given time horizon $T>0$. Each pair of vertices $u,v\in\langle N\rangle$ can be connected by an edge (which we associate with the state ‘1’) or not (which we associate with the state ‘0’). Throughout this paper, the resulting edge process of the vertex-pair $(u,v)$ is denoted by $(a_{N,u,v}(t))_{t\in[0,T]}$. We restrict ourselves to what could be called homogeneous simple graphs, i.e., across all $u,v\in\langle N\rangle$ with $u<v$ and for any fixed $N\in\mathbb{N}$, the processes $(a_{N,u,v}(t))_{t\in[0,T]}$ are independent, identically distributed, and $G_{N}(t)$ does not have any self-loops and double edges. This in particular entails that (i) $a_{N,u,u}(t)=0$ for all $N\in\mathbb{N}$, and $t\in[0,T]$, and (ii) $a_{N,u,v}(t)=a_{N,v,u}(t)\in\{0,1\}$ for all $N\in\mathbb{N}$, $u,v\in\langle N\rangle$, $u\not=v$ and $t\in[0,T]$.
@@ 2. Model
The papers on asymptotic normality for static Erdős-Rényi random graphs, such as the ones cited in the introduction, work with some assumption on the behavior of the edge probability in the regime that the number of vertices $N$ grows large. In our dynamic counterpart we have to impose a similar assumption.
@@
$\lim_{N\to\infty}N^{v}\varrho_{N}^{e}=\lim_{N\to\infty}N^{v-e\frac{v^{\star}}{e^{\star}}}=\begin{cases}0\quad&\text{if }v<e\frac{v^{\star}}{e^{\star}},\\ 1\quad&\text{if }v=e\frac{v^{\star}}{e^{\star}},\\ \infty\quad&\text{if }v>e\frac{v^{\star}}{e^{\star}},\end{cases}$
@@
[…]
Remark 2.3.
@@
[…]
One of the most natural class of processes which satisfies Assumptions 2.1 and 2.4 is the class of alternating processes $a_{N}(\cdot)$, where $a_{N}(t)\in\{0,1\}$ for all $t\in[0,T]$. An alternating process is initiated by sampling the Bernoulli random variable $a_{N}(0)$, and then it changes between on and off at random points in time. We let $\xi_{N,x,i}$, for $x\in\{0,1\}$ and $i,N\in\mathbb{N}$, denote the duration of the $i$-th ‘sojourn time’ given that process starts at the state $x$, where such a sojourn time is either an on-time or an off off-time, depending on the parity of $x+i$; it is typically assumed that all the on- and off-times are independent. If these random variables are in addition from a continuous distribution, then under a couple of mild restrictions such a process satisfies (2.5) and (2.6) by virtue of the following proposition.
Proposition 2.6.
@@
As a particular example of the alternating processes that satisfy the assumptions of Proposition 2.6, we can consider a stationary alternating renewal process with exponential on- and off-times. This concretely means that $\xi_{N,x,i}$ has an exponential distribution with parameter $\lambda_{N,{\rm off}}$ if $x+i$ is even and with parameter $\lambda_{N,{\rm on}}$ if $x+i$ is odd, and $p_{N}(0)=\lambda_{N,{\rm off}}/(\lambda_{N,{\rm on}}+\lambda_{N,{\rm off}})$. In this case we have for any $t\in[0,T]$
@@
When considering the specific parametrization $\lambda_{N,{\rm off}}=N^{-\alpha}\lambda_{{\rm off}}$ and $\lambda_{N,{\rm on}}\equiv\lambda_{{\rm on}}$ for some parameters $\lambda_{\rm on},\lambda_{\rm off},\alpha>0$, we obtain that the process $a_{N}(\cdot)$ satisfies Assumption 2.1 and 2.4 with $\varrho_{N}=N^{-\alpha}$,
@@
3. Graph notations
In this section we introduce various graph-related notations.
@@ 3. Graph notations
    In the first place, for two graphs $H$ and ${H^{\star}}$ we write ${H}\sim{H^{\star}}$ if they are isomorphic, and we write ${H}\subset{H^{\star}}$ if $H$ is a subgraph of ${H^{\star}}$. Denote by $K_{N}$ a complete graph with vertices labeled $1,\ldots,N$.
@@
    For any graph $H$ we denote by $\mathcal{V}(H)$ ($\mathcal{E}(H)$, respectively) the number of vertices (edges, respectively) in $H$; we also use the notation $\mathcal{A}(H)$ to denote the number of automorphisms of $H$.
@@
    In addition to conventional subgraphs, we define embedded subgraphs $g{\,\sqsubset\,}K_{N}$ as a subgraph of $K_{N}$ with particularly labeled vertices ${\tt V}(g)=\{v_{1},\ldots,v_{a}\}\subset\langle N\rangle$, and the edge set
    ${\tt E}(g)\subset\bigcup_{\begin{subarray}{c}i,j=1\\ i<j\end{subarray}}^{a}\{(v_{i},v_{j})\}.$
    We treat two embedded subgraphs as different if they have different sets of vertices or different sets of edges. For any pair of embedded subgraphs $g_{1},g_{2}\sqsubset K_{N}$, we can naturally define the embedded subgraphs $g_{1}\cup g_{2}{\,\sqsubset\,}K_{N}$ as a subgraph with vertices ${\tt V}(g_{1}\cup g_{2})={\tt V}(g_{1})\cup{\tt V}(g_{1})$ and edges ${\tt E}(g_{1}\cup g_{2})={\tt E}(g_{1})\cup{\tt E}(g_{2})$. The embedded graphs $g_{1}\cap g_{2}{\,\sqsubset\,}K_{N}$ and $g_{1}\setminus g_{2}$ can be defined in a similar way. We write $g_{1}{\,\sqsubset\,}g_{2}$ for two embedded subgraphs $g_{1},g_{2}{\,\sqsubset\,}K_{N}$ if ${\tt V}(g_{1})\subset{\tt V}(g_{2})$ and ${\tt E}(g_{1})\subset{\tt E}(g_{2})$. In particular, for any $t\in[0,T]$ the random graph $G_{N}(t)$ can be considered as an embedded subgraph $G_{N}(t)\,\sqsubset\,K_{N}$, so that in this way we can define the relation $g\,\sqsubset\,G_{N}(t)$ for any embedded subgraph $g\,\sqsubset\,K_{N}$.
@@
    The main object that is studied in this paper, concerns the $m$-dimensional subgraph count process $(X_{N,1}(\cdot),\ldots,X_{N,m}(\cdot))$, whose $i$-th entry represents the number of subgraphs of $G_{N}(\cdot)$ which are isomorphic to some fixed graphs $H_{i}$. Given any sequence of subgraphs $H_{1},\ldots,H_{m}$, using the fact that $G_{N}(t)$ has a natural embedding in $K_{N}$, we can interpret $X_{N,i}(t)$ for $i\in\langle m\rangle$ as the number of embedded subgraphs $g{\,\sqsubset\,}K_{N}$ such that $g\sim{H}_{i}$ and $g{\,\sqsubset\,}G_{N}(t)$. As such, it is convenient to define the set of all possible embedded subgraphs $g{\,\sqsubset\,}K_{N}$ such that they are isomorphic to some fixed graph $H$:
    ${\tt G}_{N}({H})=\{g{\,\sqsubset\,}K_{N}\colon g\sim{H}\}.$
@@
    For example, if $H$ is a triangle, then $\left\lvert{\tt G}_{N}(H)\right\rvert=\binom{N}{3}$. Define also for any ${g\,\sqsubset\,K_{N}}$ the indicator function of this subgraph $g$ being a subgraph of $G_{N}(t)$ as well:
    $\mathbb{I}_{N}(g,t)=\mathbb{I}\{g{\,\sqsubset\,}G_{N}(t)\}=\prod_{(u,v)\in{\tt E}(g)}a_{N,u,v}(t).$ (3.2)
@@
    Throughout the paper we need to control the way how different subgraphs are intersecting. So, define for any pair of graphs $H$, ${H}^{\star}$ a set of all common subgraphs:
    ${\tt CS}({H,H^{\star}})=\{g\colon g\subset{H},\,g\subset{H^{\star}}\}.$
    For brevity we write ${\tt CS}({H})={\tt CS}({H},{H})$, which basically stands for the set of all subgraphs of graph $H$. For any graph $H$ with at least one edge we define
@@
    $\lim_{N\to\infty}N^{-\mathcal{V}(g)}\varrho_{N}^{-\mathcal{E}(g)}=0.$ (3.4)
    Equivalently, Assumption 3.1 tells us that for any $g\subset H$ such that $\mathcal{V}(g)\geqslant 1$
    $\lim_{N\to\infty}\mathcal{F}_{N}(g)=0.$
@@
    Define $X_{N}(t)$ as the number of inclusions of the graph $H$ in $G_{N}(t)$, and $X_{N}^{\star}(t)$ as the number of inclusions of the subgraph $g$ in $G_{N}(t)$. Then,
@@
[…]
    The asymptotic behavior of functions $\mathcal{F}_{N}(g)$ may be different for different subgraphs $g$. So, it is useful to identify subgraphs for which the decay of $\mathcal{F}_{N}(g)$ is the slowest (i.e., subgraphs $g$ for which $\mathcal{F}_{N}(g)$ remains asymptotically at least as large as the decay $\mathcal{F}_{N}(g^{\star})$ of any other subgraph $g^{\star}$):
    ${\tt OCS}({H},{H^{\star}})=\left\{g\in{\tt CS}({H},{H^{\star})},\mathcal{E}(g)\geqslant 1\colon\begin{array}[]{l}\forall g^{\star}\in{\tt CS}({H},{H^{\star}})\,\mathcal{E}(g^{\star})\geqslant 1,\\ \lim_{N\to\infty}\mathcal{F}_{N}(g^{\star})/\mathcal{F}_{N}(g)\in[0,\infty)\end{array}\right\}.$ (3.8)
@@
    Hence, for any graphs ${H},{H^{\star}}$, if we define $\mathcal{F}^{\rm opt}_{N}({H},{H^{\star}})=\mathcal{F}_{N}(g^{\star})$ for some subgraph $g^{\star}\in{\tt OCS}(H,H^{\star})$, then due to (3.9) for any $g\in{\tt CS}({H},{H^{\star}})$, $\mathcal{E}(g)\geqslant 1$,
@@
    Again, for simplicity we write $\mathcal{F}^{\rm opt}_{N}({H}):=\mathcal{F}^{\rm opt}_{N}({H},{H})$. We call two graphs $H$ and ${H^{\star}}$ equioptimal if
@@
    Again, by virtue to (2.4), the limit in (3.11) cannot be positive, finite and different from $1$$\Diamond$}
@@
    We write ${H}\cong{H^{\star}}$ if graphs $H$ and ${H^{\star}}$ are equioptimal. A combination of (3.11) and (3.10) tells us that if ${H}\cong{H^{\star}}$ then

### v2

# for the simultaneous subgraph count of

@@
[…]
for the simultaneous subgraph count of
[…]
1 Introduction
Over the past decades, inspired by applications in a broad range of disciplines (such as biology, physics, economics, and computer science), various types of random graph models have been introduced and analyzed. Random graph models are specified via an underlying probabilistic mechanism. The Erdős–Rényi random graph [7, 11], in which there are $N$ vertices and in which every possible vertex pair is connected independently with a given probability $p\in(0,1)$, is arguably the most well-known.
Most literature on random graphs focuses on static models, which represent a single instance of the random graph without accounting for them changing over time. However, real-world networks typically do evolve over time, motivating the study of dynamic random graphs, particularly dynamic Erdős–Rényi random graphs (e.g., [25, 20, 18, 17, 5]). In these models, edges evolve independently over time, alternating between being present and absent. Essentially all properties that have been studied for static random graphs can be analyzed for their dynamic counterparts as well. For instance, in the context of dynamic random graphs there is now a notion of a giant component process, and a notion of a spectrum process. In the present paper we concentrate on the subgraph counts process, in that our aim is to describe the joint evolution of the entries of a subgraph count vector. We thus capture, for instance, the joint distribution of the number of edges and the number of triangles over time.
@@ 1 Introduction
The first results showing asymptotic normality for subgraph counts in a static Erdős–Rényi random graph appeared already in [6]: an appropriately centered and normalized version of the subgraph count converges, as $N\to\infty$, to a zero-mean Gaussian random variable. These results were further generalized to the regime where the edge probability $p=p_{N}$ tends to zero as the number of vertices $N$ grows large; see e.g. [23] that covers specific types of subgraphs. Finally, in [24] necessary and sufficient conditions on the decay of $p_{N}$ were established for obtaining asymptotic normality for any subgraph count.
@@
Importantly, there seem to be no asymptotic normality results covering the setting of random graphs that evolve over time, even in the simplest case in which the edge processes do not depend on the number of vertices $N$. The main objective of this paper is to settle this, by establishing a joint functional central limit theorem for dynamic Erdős–Rényi random graphs. Concretely, we identify conditions under which an appropriately centered and normalized version of a vector of subgraph counts has, at the process level, a Gaussian limit. Moreover, we explicitly characterize the limiting Gaussian process. The type of convergence considered is in distribution, in $\mathbb{D}([0,T],\mathbb{R}^{m})$, which is the set of all càdlàg functions $f(\cdot)\colon[0,T]\to\mathbb{R}^{m}$, equipped with $J_{1}$ distance, i.e., for any functions $f(\cdot),\,g(\cdot)\in\mathbb{D}([0,T],\mathbb{R}^{m})$
$d\bigl(f(\cdot),\,g(\cdot)\bigr)=\inf_{\lambda\in\Lambda}\max\left(\sup_{t\in[0,T]}\left\lvert\lambda(t)-t\right\rvert,\,\sup_{t\in[0,T]}\|f(t)-g(\lambda(t))\|\right),$
for some finite, deterministic horizon $T$, where $\Lambda$ is the class of strictly increasing, continuous mappings of $[0,T]$ onto itself. We refer to e.g. [3, Section 12] for a more detailed account.
In our setup, we do not focus on one specific type of edge dynamics, in that we aim to work with a specification of the dynamic Erdős–Rényi random graph that is as general as possible. We follow the ‘traditional’ path in proving process-level convergence: we first establish finite-dimensional convergence, followed by a tightness step that yields the desired functional convergence. In order to accommodate this tightness, we have to impose some assumptions on the edge dynamics. More concretely, in our framework the probabilities of the edge switching once or twice (in an interval of given length, that is) are assumed to obey a Lipschitz-type condition. At the methodological level, some of the steps of our analysis follow their counterparts in [13], where the principal eigenvalue process of the graph’s adjacency matrix is analyzed in the regime in which the edge processes do not depend on $N$.
The main result of this paper, Theorem 4.1, establishes that the limiting behavior (as the size of our graph grows large) of the subgraph count process is Gaussian, and strictly correlated with the subgraph count for the least common subgraph of our graph. This least common subgraph is defined as a subgraph with at least one edge and the asymptotically smallest expected number in our dynamic Erdős–Rényi random graph. For instance, if the edge process does not depend on the size of the random graph (as before denoted by $N$), the expected subgraph count asymptotically scales as $N^{v}$, where $v$ is the size of the subgraph under consideration. In this scenario, the least common subgraph for any graph with at least one edge is a single edge, and thus every subgraph count process is strictly correlated with the edge count. This implies that all possible subgraph counts are strictly correlated with one another. We provide a detailed exploration of this phenomenon in Corollary 4.2.
@@
2 Model
We start by constructing our dynamic Erdős–Rényi random graph process, in the sequel denoted by $G_{N}(t)$, with $N\in\mathbb{N}$ vertices over the time interval $t\in[0,T]$ for some given time horizon $T>0$. Each pair of vertices $u,v\in\langle N\rangle$ can be connected by an edge (which we associate with the state ‘1’) or not (which we associate with the state ‘0’). Throughout this paper, the resulting edge process of the vertex-pair $(u,v)$ is denoted by $(a_{N,u,v}(t))_{t\in[0,T]}$. We restrict ourselves to what could be called homogeneous simple graphs, i.e., across all $u,v\in\langle N\rangle$ with $u<v$ and for any fixed $N\in\mathbb{N}$, the processes $(a_{N,u,v}(t))_{t\in[0,T]}$ are independent, identically distributed, and $G_{N}(t)$ does not have any self-loops and double edges. This in particular entails that (i) $a_{N,u,u}(t)=0$ for all $N\in\mathbb{N}$, and $t\in[0,T]$, and (ii) $a_{N,u,v}(t)=a_{N,v,u}(t)\in\{0,1\}$ for all $N\in\mathbb{N}$, $\in\langle N\rangle, uvu\not=v$ and $t\in[0,T]$.
@@ 2 Model
The papers on asymptotic normality for static Erdős–Rényi random graphs, such as the ones cited in the introduction, work with some assumption on the behavior of the edge probability in the regime that the number of vertices $N$ grows large. In our dynamic counterpart we have to impose a similar assumption.
@@
${\lim_{N\to\infty}N^{v}\varrho_{N}^{e}=\lim_{N\to\infty}N^{v-e\frac{v^{\star}}{e^{\star}}}=\begin{cases}0\quad&\text{if }e/v>e^{\star}/{v^{\star}},\\ 1\quad&\text{if }e/v=e^{\star}/{v^{\star}},\\ \infty\quad&\text{if }e/v<e^{\star}/{v^{\star}}.\end{cases}}$
@@
Technically, to check (2.3) we need to have $v,e\in\mathbb{N}$, but at various places in our proofs we work with powers that potentially have a negative sign; this explains why in (2.3) we chose to include $v,e\in\mathbb{Z}$. $\Diamond$}
Remark $2.$
@@
for all $0\leqslant r\leqslant s\leqslant T$. It implies that the function $p_{N}(\cdot)/\varrho_{N}$ is continuous for any $N\in\mathbb{N}$. Hence, due to (2.1), there exists constant $p_{+}\in(1,\infty)$ such that
${\sup_{N\in\mathbb{N}}\sup_{t\in[0,T]}}\frac{p_{N}(t)}{\varrho_{N}}$ $\leqslant p_{+},$ (2.7)
$\Diamond$}
One of the most natural class of processes which satisfies Assumptions 2.1 and 2.2 is the class of alternating processes $a_{N}(\cdot)$, where $a_{N}(t)\in\{0,1\}$ for all $t\in[0,T]$. An alternating process is initiated by sampling the Bernoulli random variable $a_{N}(0)$, and then it changes between on and off at random points in time. We let $\xi_{N,x,i}$, for $x\in\{0,1\}$ and $i,N\in\mathbb{N}$, denote the duration of the $i$-th ‘sojourn time’ given that process starts at the state $x$, where such a sojourn time is either an on-time or an off-time, depending on the parity of $x+i$; it is typically assumed that all the on- and off-times are independent. If these random variables are in addition from a continuous distribution, then under a couple of mild restrictions such a process satisfies (2.4) and (2.5) by virtue of the following proposition.
Proposition 1.
@@
As a particular example of the alternating processes that satisfy the assumptions of Proposition 1, we can consider a stationary alternating renewal process with exponential on- and off-times. This concretely means that $\xi_{N,x,i}$ has an exponential distribution with parameter $\lambda_{N,{\rm off}}$ if $x+i$ is even and with parameter $\lambda_{N,{\rm on}}$ if $x+i$ is odd, and $p_{N}(0)=\lambda_{N,{\rm off}}/(\lambda_{N,{\rm on}}+\lambda_{N,{\rm off}})$. In this case we have for any $t\in[0,T]$
@@
When considering the specific parametrization $\lambda_{N,{\rm off}}=N^{-\alpha}\lambda_{{\rm off}}$ and $\lambda_{N,{\rm on}}\equiv\lambda_{{\rm on}}$ for some parameters $\lambda_{\rm on},\lambda_{\rm off},\alpha>0$, we obtain that the process $a_{N}(\cdot)$ satisfies Assumptions 2.1 and 2.2 with $\varrho_{N}=N^{-\alpha}$,
@@
3 Graph notations
In this section we introduce various graph-related notations. Due to the dynamic nature of our graph process, we need to introduce some new concepts that were not required in the static setting.
@@ 3 Graph notations
    Throughout the paper we consider only non-empty graphs whose vertex set is a subset of $\mathbb{N}$. We denote the set of such graphs by $\tt G$$_{\infty}$. For any graph $H\in{\tt G}_{\infty}$ we denote by ${\tt V}(H)$ (${\tt E}(H)$, respectively) the set of vertices (edges, respectively) of the graph $H$. By definition of the set $\tt G$$_{\infty}$, ${\tt V}(H)\subset\mathbb{N}$ and ${\tt E}(H)\subset\{(u,v)\in\mathbb{N}^{2}\colon u<v\}$. In particular, for any $t\in[0,T]$ the random graph $G_{N}(t)$ can be considered as a random element of $\tt G$$_{\infty}$ with NN vertices.
@@
    For two graphs $H,{H^{\star}}\in{\tt G}_{\infty}$ we write ${H}\sim{H^{\star}}$ if they are isomorphic, and we write ${H}\subset{H^{\star}}$ if $H$ is a subgraph of $H^{\star}$}, i.e., ${\tt V}(H)\subseteq{\tt V}(H^{\star})$ and ${\tt E}(H)\subseteq{\tt E}(H^{\star})$. Denote for any $N\in\mathbb{N}$ by $K_{N}$ a complete graph with vertices labeled $1,\ldots,N$, i.e., ${\tt V}(K_{N})=\langle N\rangle$ and ${\tt E}(K_{N})=\{(u,v)\in\mathbb{N}^{2}\colon u<v\}$.
@@
    For any graph $H\in{\tt G}_{\infty}$ we denote by $\mathcal{V}(H)$ ($\mathcal{E}(H)$, respectively) the number of vertices (edges, respectively) in $H$, i.e., $\mathcal{V}(H)=\left\lvert{\tt V}(H)\right\rvert$ and $\mathcal{E}(H)=\left\lvert{\tt E}(H)\right\rvert$; we also use the notation $\mathcal{A}(H)$ to denote the number of automorphisms of $H$.
@@
    For any pair $g_{1},g_{2}\in{\tt G}_{\infty}$ we use the notation $g_{1}\cup g_{2}$ for the graph with ${\tt V}(g_{1}\cup g_{2})={\tt V}(g_{1})\cup{\tt V}(g_{1})$ and ${\tt E}(g_{1}\cup g_{2})={\tt E}(g_{1})\cup{\tt E}(g_{2})$. The graphs $g_{1}\cap g_{2}$ and $g_{1}\setminus g_{2}$ can be defined in a similar way.
  • $\circ$
    The main object that is studied in this paper, concerns the $m$-dimensional subgraph count process $(X_{N,1}(\cdot),\ldots,X_{N,m}(\cdot))$, whose $i$-th entry represents the number of subgraphs of $G_{N}(\cdot)$ which are isomorphic to some fixed graphs $H_{i}\in{\tt G_{\infty}}$. Given any sequence of subgraphs $H_{1},\ldots,H_{m}{\in{\tt G}_{\infty}}$, we can interpret $X_{N,i}(t)$ for $i\in\langle m\rangle$ as the number of graphs $g\,{\in{\tt G}_{\infty}}$ such that $g\sim{H}_{i}$ and $g{\,{\subset}\,}G_{N}(t)$. Using that $G_{N}(t)\subset K_{N}$ for any fixed $N\in N$ almost surely, we can consider only such graphs $g\,\in{\tt G}_{\infty}$ that ${\tt V}(g)\subseteq\langle N\rangle$. As such, it is convenient for any $H\in{\tt G}_{\infty}$ to define the set of all possible graphs $g\,{\in{\tt G}_{\infty}}$ with ${\tt V}(g)\subseteq\langle N\rangle$ such that they are isomorphic to some fixed graph $H$:
    ${\tt G}_{N}({H})={\{g\in{\tt G}_{\infty}\colon{\tt V}(g)\subseteq\langle N\rangle,\,g\sim{H}\}}.$
@@
    For example, if $H$ is a triangle, then $\left\lvert{\tt G}_{N}(H)\right\rvert=\binom{N}{3}$. Define also for any $g\,{\in{\tt G}_{\infty}}$} the indicator function of this graph $g$ being a subgraph of $G_{N}(t)$:
    $\mathbb{I}_{N}(g,t)=\mathbb{I}\{g{\,{\subset}\,}G_{N}(t)\}=\prod_{(u,v)\in{\tt E}(g)}a_{N,u,v}(t).$ (3.2)
@@
    Throughout the paper we need to control the way how different subgraph patterns can be obtained when the graphs are intersecting. To formalize this concept, denote the quotient set of all graphs with at least one edge by ${\tt G}_{\sim}=\{g\in{\tt G}_{\infty}\colon\mathcal{E}(g)\geqslant 1\}/\sim$, where $\sim$ was introduced above. We refer to any element of $\tt G$$_{\sim}$ as a subgraph pattern. We denote by $\pi:{\tt G}_{\infty}\to{\tt G}_{\sim}$ the classical projection on the quotient set. Then, define for any pair of graphs $H,\,{H}^{\star}\in{\tt G}_{\infty}$ a set of all common subgraph patterns with at least one edge:
    ${\tt CS}({H,H^{\star}})=\{g\in{\tt G}_{\sim}\colon\exists g_{1},g_{2}\in{\tt G}_{\infty}\colon g_{1}\subset{H},\,g_{2}\subset{H^{\star}},\,\pi(g_{1})=\pi(g_{2})=g\}.$
    A graph pattern $g\in{\tt CS}({H,H^{\star}})$ represents an unlabeled subgraph structure common to both $H$ and $H^{\star}$.
    For example, let $H$ be a triangle with vertex set $\{1,2,3\}$ and $H^{\star}$ an edge with vertices $\{4,5\}$. Then, even though graphs $H$ and $H^{\star}$ do not have any common subgraph, they still have a common subgraph pattern, which is the edge pattern.
    For brevity we write ${\tt CS}({H})={\tt CS}({H},{H})$, which basically stands for the set of all subgraph patterns of graph $H$. Note that functions $\mathcal{V}(\cdot)$, $\mathcal{E}(\cdot)$ and $\mathcal{A}(\cdot)$ are invariant with respect to graph isomorphism, so we can as well define them for the subgraph patterns.
    In addition, for any graph $H\in{\tt G}_{\infty}$ and any graph pattern $g\in{\tt G}_{\sim}$ we define by $\mathcal{S}(H,g)$ the total number of different subgraphs of $H$ which are isomorphic to $g$, i.e.,
    $\mathcal{S}(H,g)=\sum_{g_{0}\,\subset\,{H}}\mathbb{I}\{\pi(g_{0})=g\}.$ (3.3)
  • $\circ$
    For any graph $H{\in{\tt G}_{\infty}}$ (or graph pattern $H\in{\tt G}_{\sim}$) with at least one edge we define for any $N\in\mathbb{N}$
@@
    $\lim_{N\to\infty}{\mathcal{F}_{N}(g)}=0.$ (3.5)
@@
    Denote by $X_{N}(t)$ the number of subgraphs of $G_{N}(t)$ which are isomorphic to $H$, and $X_{N}^{\star}(t)$ the number of subgraphs of $G_{N}(t)$ which are isomorphic to $g$.
    Then,
@@
    The conclusion of the above considerations is that Assumption 3.1 is strictly necessary to ensure a non-degenerate subgraph count as $N\to\infty$. $\Diamond$}
    The asymptotic behavior of functions $\mathcal{F}_{N}(g)$ may be different for different subgraphs $g$. It is therefore useful to identify, for any pair of graphs $H,H^{\star}\in{\tt G}_{\infty}$, common subgraph patterns for which the decay of $\mathcal{F}_{N}(g)$ is the slowest. More, concretely, we consider common subgraph patterns $g$ for which $\mathcal{F}_{N}(g)$ remains asymptotically at least as large as the decay $\mathcal{F}_{N}(g^{\star})$ of any other common subgraph pattern $g^{\star}$:
    ${{\tt OCS}({H},{H^{\star}})=\left\{g\in{\tt CS}({H},{H^{\star})}\colon\forall g^{\star}\in{\tt CS}({H},{H^{\star}}):\lim_{N\to\infty}\frac{\mathcal{F}_{N}(g^{\star})}{\mathcal{F}_{N}(g)}\in[0,\infty)\right\}.}$ (3.7)
@@
    Hence, for any graphs ${H},{H^{\star}}\in{\tt G}_{\infty}$, if we define $\mathcal{F}^{\rm opt}_{N}({H},{H^{\star}})=\mathcal{F}_{N}(g^{\star})$ for some $g^{\star}\in{\tt OCS}(H,H^{\star})$, then due to (3.8) for any $g\in{\tt CS}({H},{H^{\star}})$,
@@
    Again, for simplicity we write $\mathcal{F}^{\rm opt}_{N}({H}):=\mathcal{F}^{\rm opt}_{N}({H},{H})$. We call two graphs $H,\,H^{\star}\in{\tt G}_{\infty}$ equioptimal if
@@
    Again, by virtue to (2.3), the limit in (3.10) cannot be positive, finite and different from 11. $\Diamond$}
@@
    We write ${H}\cong{H^{\star}}$ if graphs $H,\,{H^{\star}}\in{\tt G}_{\infty}$ are equioptimal. A combination of (3.10) and (3.9) tells us that if ${H}\cong{H^{\star}}$ then